Points (2,4) and (5,__) lie on a line whose slope is 7/3.
Answer by mananth(16949)

Points (2,4) and (5,__) lie on a line whose slope is 7/3.
slope = m = (y2-y1)/(x2-x1)
(x-4)/(5-2)= 7/3
(x-4) = 3*7/3
x-4=7
x=11
..
(5,11)